OpenStudy (anonymous):

multiply and simply by factoring ^4√10,000 * ^4√100,000 =

OpenStudy (anonymous):

\[\sqrt[4]{10^4}\times \sqrt[4]{10^5}=\sqrt[4]{10^9}\] \[=100\sqrt[4]{10}\]